A 1100 kg car pulls a boat on a trailer. (Enter the magnitude.) 


(a) What total force (in N) resists the motion of the car, boat, and trailer, if the car exerts a 1930-N force on the road and produces an acceleration of 0.530 m/s2? The mass of the boat plus trailer is 620 kg. 


(b) What is the force (in N) in the hitch between the car and the trailer if 80% of the resisting forces are experienced by the boat and trailer ?
Answer #1

4. (a) Total resistance force, Fr

Fr = Fc - M*a

Where, M is the total mass of the car, boater, trailer

M = 1100 + 620 = 1720 kg

a is the acceleration = 0.530 m/s^2

Fc, force exerted by the car on the road = 1930 N

Fr = 1930 - (1720*0.530)

= 1018.4 N

(b) 80% of total resistance force experienced by boat and trailer = = (80/100) * 1018.4 = 814.72 N

The net force needed to accelerate = 620*0.53 = 328.6 N

The total force needed to accelerate the boat and trailer,

= 814.72 + 328.6 N = 1143.32 N

The force hitched between the car and the trailer, = 1143.32 N